• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name fieschi

    Log entry time 00:08:28 on December 1,2004

    Entry number 137573

    keyword=shift summary

    19.50 HV calo swithed ON

    20.15 problem with moller -> restore running configuration
    target moved to hydrogen (15 cm)

    20.20 harp scan -> OK

    20.30 start elastic calibration run 4994 (4uA)
    22.30 bleedthrough measurement
    23.10 move Left HRS to 37.69 degree
    momentum set to 2.381 GeV/c

    elastic calibration
    __________________________________
    LHRS angle 36.48 degree, momentum : 2.489 GeV/c
    DVCS angle -20.50 degree
    ___________________________________
    4994
    4995
    ____________________________________
    LHRS angle 37.69 degree, momentum : 2.381 Gev/c
    --------------------------------------
    4996 running


    Beam trip
    ________________________________________
    20.00 -> 21.00 8*20s -> 3min
    21.00 -> 22.00 18*20s -> 6min
    22.00 -> 23.00 18*20s -> 6min
    23.00 -> 00.00 24*20s -> 8min


    compton
    ______________________
    run pol
    9748 74.3 %